Where



C=250+0.75Yd



I=350



G=800



X=500



M=200+0.4Y



T=0.15Y




Use the above information to answer the following questions




1) if C=250+0.75Yd interpret the meaning of the value 250



2)state the values of MPT,MPC and MPS on disposable income and MPM


1
Expert's answer
2022-01-25T08:45:12-0500

1 If C=250+0.75Yd, the meaning of the value 250 is autonomous consumption. Autonomous consumption can be defined as the consumption of consumers even when they have no disposable income. This kind of consumption can be finance through loans, the consumer's previous savings, etc.

Mathematically, C=a+bY

Where;

C= consumption

a= autonomous consumption

b= marginal propensity to consume

Y= disposable income

The autonomous consumption is the consumption when disposable income equals zero.

Substituting disposable income with zero in the given function we have,

C=250+0.75(0)

C=250+0

Therefore, C=250 which is the autonomous consumption.

2 (a) MPT - Marginal Propensity to Tax is the ratio of a change in Tax (T) to a change in income (Y).

"\text{MPT}=\dfrac{\Delta T}{\Delta Y}\\ 
\text{T}=0.15Y\\
Therefore, MPT=0.15."

(b) MPC - Marginal Propensity to Consume is the ratio of a change in consumption to a change in income.

"MPC=\dfrac{\Delta C}{\Delta Y}\\
C= 250+0.75Y_d\\
Therefore, MPC=0.75"

(c) MPS = Marginal Propensity to Save

Since MPC+MPS=1,

We can say that MPS=1-MPC

Substituting for MPC, we have,

MPS=1-0.75=0.25

Therefore, MPS=0.25.

(d) MPM - Marginal Propensity to Import is the ratio of a change in import to a change in income.

"MPM=\dfrac{\Delta M}{\Delta Y}\\
M=200+0.4Y
Therefore, MPM=0.4"